OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 8m and a depth of 1.6m . If the water is pumped out of the pool at the rate of 11m^3 per hour, how many hours does it take to empty the pool? Use the value 3.14 for Pi , and round your answer to the nearest hour.

OpenStudy (anonymous):

V = 4^2 pi * 1.6 = 80.4 m^3 -> 80.4 / 11 = 7 hrs

OpenStudy (anonymous):

why so many medal Faman

OpenStudy (anonymous):

I'm checking whether I'm deserved for the medals :P

OpenStudy (anonymous):

Thank you @KendrickLamar2014

OpenStudy (kendricklamar2014):

No problem @Beautykillzz